OpenStudy (anonymous):

The Tacoma Tigers were in last place in their division with a record of 29 wins and 43 losses. In other words, their current winning percentage was 29/29+43=29/72=.403. We are going to find out how many games in a row the Tigers have to win to increase their winning percentage to .500. To set up an equation, consider this: The number of games won in a row would have to be added to both the number of wins and the total number of games, and the result would have to be .500. What equation represents this situation?

OpenStudy (zarkon):

\[\frac{n+29}{n+29+43}=.5\]

OpenStudy (anonymous):

29+x/43+x=.5?

OpenStudy (zarkon):

no

OpenStudy (anonymous):

It's a multiple choice question. So yours would basically be 29+x/72+x=.5. Thanks!

OpenStudy (zarkon):

yes